OpenStudy (anonymous):

Help please! Find the x-intercept of -2x + y = 4?

OpenStudy (alexandervonhumboldt2):

substitute 0 for y and solve

OpenStudy (alexandervonhumboldt2):

-2x+0=4 -2x=4 x=-2

OpenStudy (lochana):

for what value of y do you think it intercepts x axis?

OpenStudy (lochana):

yeah. that is what I about to say:)